Determine the primary and reciprocal trig ratios for the following angles.


a) (7pi)/4  


b) (- 2pi)/3


1
Expert's answer
2021-02-16T09:57:39-0500

Primary:

"sin(7\pi/4) = -(\sqrt2/2)"

"cos(7\pi/4) = \sqrt2/2"

"tan(7\pi/4) = -1"


"sin(-2\pi/3) = -(\sqrt3/2)"

"cos(-2\pi/3) = -(1/2)"

"tan(-2\pi/3) = \sqrt3"


Reciprocal:

"sec(7\pi/4) = \sqrt2"

"csc(7\pi/4) = -\sqrt2"

"cot(7\pi/4) = -1"


"sec(-2\pi/3) = -2"

"csc(-2\pi/3) = -(2/\sqrt3)"

"cot(-2\pi/3) = \sqrt3/3"